# A kinetic model of solar wind acceleration driven by ambipolar electric potential and velocity-space diffusion

### v2

@@ 3 Exospheric approach on diffusion
The main limitation of ordinary kinetic exospheric models of the solar wind is their inability to self-consistently define a halo population outside the regions of velocity space described in the previous section. To solve this problem, we introduce a halo population that is directly dependent on the number of particles in the classically filled regions of velocity space (A, B, and C). In this way, the radial evolution of the total density stays unchanged while the electrons are redistributed in velocity space so that a tenuous halo appears for greater perpendicular velocities.
Therefore, diffusion is modeled by defining a new parameter that aims to reduce the number of electrons in the A, B, and C regions and adding the same number of electrons to the initially empty region of Fig. 2 labeled as D. This diffusion parameter is defined as
@@
so that when $r\rightarrow\infty$ ($\eta\rightarrow 0$) then $D_{r}\rightarrow\zeta$, with $\zeta$ being a value between 0 and 1 that modulates the relative number of diffused particles, compared to the others (of regions A, B, and C), at an infinite distance from the Sun. The choice that the diffusion profile evolves as $1-(r_{0}/r)^{6}$ is entirely arbitrary. It was selected to ensure that $D_{r}=0$ at $r=r_{0}$ while also preventing too rapid growth, thereby avoiding the diffusion coefficient reaching near its maximum value well before approaching 1 AU.
Furthermore, the density within the combined velocity space region of A, B, and C can be deduced by the addition of Equations (5), (25), and (30) of Pierrard & Lemaire (1996). This can be written as
@@
so that the density of electrons $n_{0}$ is equal to the total density of electrons at the exobase since
@@
Using the total density and the diffusion parameter, we can now deduce the density of the depleted region (combined region of A, B, and C) which corresponds to
@@
so that the total density remains unchanged since $n_{e}(r)=n_{e}^{\prime}(r)+n_{e,\textbf{D}}(r)$.
The electron flux is here defined using Equation (26) of Pierrard & Lemaire (1996) by taking into account the diffusion. This corresponds to the only nonzero contribution of the flux associated with the escaping particles (region B) and written as
@@
From which it follows that the total flux is $F_{e}^{\prime}(r)+F_{e,\textbf{D}}(r)=F_{e}(r)(1-D_{r})$, with $F_{e,\textbf{D}}(r)=0$ being the diffusion contribution to the flux and $F_{e}(r)$ being the total flux as if there were no diffusion. Here, the diffused electrons (from the depleted region) are scattered so that they do not contribute to the flux ($F_{e,\textbf{D}}(r)=0$). This provides a simple approach for formulating the effect of diffusion on the flux in our kinetic exospheric model. Note that the proton flux must be modified accordingly to maintain the zero-current condition that was used to find the solution for the electric potential.
Figure 2: Radial evolution of the electron VDF from the exobase (1.2 $R_{s}$ in our model) to 1 AU ($\sim$215 $R_{s}$) modeled by including diffusion for electrons. The region labeled as D represents the electron diffusion region where $\zeta=0.1$ was assumed.
To deduce the electron pressures of the D region, we first define the pressure without any diffusion of electrons by adding Equations (7), (27), and (31) of Pierrard & Lemaire (1996) for the parallel component to obtain
@@
and adding Equations (8), (28), and (32) for the perpendicular component to obtain
@@
To determine the pressures in the diffusion region, we introduce the electron density $n_{e,\text{I}}$, obtained by integrating the eVDF from Equation (9) within the boundaries of the D region. This quantity corresponds to subtracting Equation (35) with (25) of Pierrard & Lemaire (1996), as expressed by
@@
In the same way, we define the parallel and perpendicular pressure by subtracting Equation (37) with (27) and Equation (38) with (28) respectively, and multiplying them by the ratio $n_{e,\textbf{D}}(r)/n_{e,\text{I}}(r)$ to obtain
@@
The pressures are those corresponding to the D region of the electron VDF in Fig. 2, where $N_{e,\textbf{D}}$ is the normalization factor of the electron VDF (replacing $N_{e}$ in Equation (9)) so that the density of this region is always equal to $n_{e,\textbf{D}}(r)$ as described by Equation (31). This is inferred by noticing that
@@
This allows us to then deduce the bulk velocity
$v_{\text{sw}}(r)\equiv\frac{F_{e}^{\prime\prime}(r)}{n_{e}^{\prime\prime}(r)}$ (47)
and total temperatures
@@ 4 Diffusive kinetic exospheric model: results
To emphasize the impact of diffusion, we analyze two scenarios: one that incorporates diffusion and another that does not. We adopt the same fundamental parameters of the exospheric model as those used for the illustrations in Figs. 1 and 2, as specified in Table 1, and we assume the diffusion radial evolution described by Equation (20) with $\zeta=0.1$ as the relative maximum amount of diffused particles. This $\zeta$ value was arbitrarily chosen to qualitatively analyze the effect of diffusion, as further improvements to the exospheric diffusion framework would be required to accurately constrain the radial evolution of diffused particles with observations. Fig. 3 shows the radial evolution of the diffusion (panel a), the electron temperatures (panel b), the solar wind speed (panel c) and the electron temperature anisotropy (panel d). These are given by our kinetic exospheric model with solid and dashed lines corresponding, respectively, to the case with ($\zeta=0.1$) and without diffusion ($\zeta=0$).
Figure 4: Radial evolution of the different electron population densities computed with the parameters of Table 1 and $\zeta=0.1$ in our modified kinetic exospheric model.
In Fig. 3a, the proportion of diffused particles increases with the radial distance, as expected from Equation (20). The radial evolution of the electron temperatures, when affected by diffusion, shows a slower decrease in Fig. 3b. This behavior is explained by the fact that the eVDF is essentially broadened by placing particles in the diffusion region. Fig. 3d shows a clear stabilization of the temperature anisotropy, which reaches a value of 1.65 at 215 $R_{s}$ when diffusion is considered. At 1 AU, observations show that the temperature anisotropy $T_{\parallel}/T_{\perp}$ is typically around 1 (that is, without anisotropy) for a slow wind of 300 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ and increases with the solar wind speed to 1.4 at 700 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ (Salem et al. 2023). In this respect, the naive introduction of diffusion in our model already brings the temperature anisotropy to a value fairly close to the observations in comparison to what was previously found with exospheric models.
The solar wind speed in Fig. 3c shows a slight decrease when diffusion is included. This is because the diffused particles are assumed to be redistributed without any preferential parallel direction. Therefore, those particles do not contribute to the bulk velocity, while they effectively diminish the number of particles within the strahl (essentially the B region), being the only contribution to the flux of the solar wind in our model. This in turn reduces the bulk velocity shown in Fig. 3c. As a consequence, because of the zero-current condition, the bulk velocity of protons must decrease by the same magnitude as that of the electrons.
@@ 5 Exospheric perspective on solar wind acceleration
As presented in Section 2, a key advantage of the kinetic exospheric model is its ability to self-consistently determine the electric field under the constraints of quasi-neutrality and zero-current conditions. Consequently, the ambipolar electric field arises directly from the greater influence of the gravitational field on protons compared to that of electrons. This resulting global electric potential contributes, at least partially, to the acceleration of the solar wind and is generally entirely responsible for the acceleration in the context of kinetic exospheric models (Berčič et al. 2021a, b).
@@
Following the kinetic exospheric approach, a signature of the ambipolar electric potential can be identified in the solar wind data. Specifically, it can be inferred from the cutoff velocity associated with the deficit in the sunward direction by analyzing the observed electron VDFs in the solar wind (Halekas et al. 2020; Berčič et al. 2021b, c; Halekas et al. 2021, 2022). Introducing the asymptotic wind speed, Halekas et al. (2022) binned the cutoff velocities of PSP between 13.3 and 50 $R_{s}$ (for encounters 3 to 11). In this way, the radial evolution of the electric potential associated with asymptotic speeds of 200–250 and 600–650 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ has been deduced and represented in Fig. 5 with dashed lines as PSP SSW (Slow Solar Wind, in green) and PSP FSW (Fast Solar Wind, in orange) respectively. An anticorrelation between the electric potential and the velocity of the solar wind can be clearly identified (Halekas et al. 2022).
@@
Figure 5: Radial evolution of the electric potential with respect to infinity. The solid lines are the electric potential retrieved by the exospheric model with the input parameters of Table 2. The corresponding solar wind terminal velocities of our model are approximately: 320 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for SSW1, 280 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for SSW2, 640 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for FSW1, and 520 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for FSW2. The dashed lines correspond to the slow solar wind (blue) and fast solar wind (orange) limits fitted to Parker Solar Probe data between 13.3 and 50 $R_{s}$ (Encounter 3 to 11) by Halekas et al. (2022). The blue dashed line is the fit of solar wind data with asymptotic speeds of 200–250 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ while the orange dashed line correspond to asymptotic speeds of 600–650 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$.
@@
For the kinetic exospheric model, the radial evolution of the self-consistent electric potential is shown in Fig. 5 for four different sets of parameters given by Table 2. The SSW1 solid black line is a slow solar wind solution found using the same test parameters as in Table 1, which was introduced to illustrate the effect of diffusion. Note that the electric potential is independent of the diffusion, since it is found by using the quasi-neutrality with an unchanged total density. It is commonly understood that exospheric models predict a faster solar wind accompanied by a larger electric potential at any radial distance, to account for the enhanced acceleration of the solar wind. This contradiction with the observed anticorrelation mentioned above holds true only if we consider that the acceleration is entirely determined by the $\kappa$ parameter. As illustrated in Fig. 5, this is evident from the comparison between SSW1 and FSW1, where reducing $\kappa$ from 6 to 3 results in a corresponding increase in the terminal wind velocity from 320 to 640 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$. However, when considering the same change in $\kappa$ for SSW2 and FSW2, a carefully selected combination of exobase levels and initial temperatures (see Table 2) produces a crossing point between the two electric potentials (at around 13 $R_{s}$) while still maintaining the increase in terminal wind velocity from 280 (for SSW2) to 520 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ (for FSW2). These parameters were chosen because they approximate the observed behavior to some extent, producing a crossing point at low altitude while maintaining realistic values (as described in the next paragraph). However, a wide range of parameter sets can reproduce a crossing point, provided they meet two key conditions: (1) a sufficiently large difference between proton and electron temperatures (compared to the difference in kappa values) in fast wind relative to slow wind, and (2) a significantly higher exobase level for slow wind. In this way, the anticorrelation between the electric potential and the solar wind terminal velocity is retrieved above 13 $R_{s}$ with the kinetic exospheric model. This indicates that the kinetic exospheric model may be better suited than previously thought for explaining the fast solar wind through the electric field, especially if the crossing point between the electric potentials of fast and slow winds occurs below 13 $R_{s}$. Although PSP observations do not appear to indicate the presence of such an intersection, or at least not above 13 $R_{s}$, the consistency of our modeled electric potentials compared with observations suggests that further investigation of the most recent PSP data is needed to completely rule out this possibility.
@@
This consideration is reinforced by the fact that the parameters found to reproduce the anticorrelation between the electric potential and the solar wind speed (see SSW2 and FSW2 of Table 2) are also consistent with the deduced coronal temperatures (David et al. 1998; Cranmer 2002). The fast solar wind typically appears above coronal holes that are characterized by a higher proton temperature compared to that of electrons and a low density that lowers the exobase compared to other regions of the Sun’s surface (Pierrard et al. 2023), which is the case for FSW2. This is also supported by the anticorrelation between the strahl parallel temperature and the solar wind speed, consistently found by Helios (65–215 $R_{s}$) and PSP (13.3–50 $R_{s}$) (Berčič et al. 2020; Halekas et al. 2022). Furthermore, as identified with SSW2, the slow solar wind is commonly associated with denser and less active regions of the equatorial corona, where the exobase is therefore higher with a generally hotter electron temperature than the coronal hole conditions (David et al. 1998). As reported by Berčič et al. 2020, these findings suggest that the strahl parallel temperature preserves the electron coronal temperature. The exospheric approach is consistent with this idea by maintaining the strahl parallel temperature constant throughout the solar wind expansion, since no collisions nor wave-particle interactions are assumed.
@@
It is important to note that the implementation of the diffusion mechanism does not inherently require that particles in region D do not contribute to the flux. For that reason, an additional acceleration mechanism could be taken into account within the current diffusion framework by considering that the transfer of core particles (combined ballistic and trapped particles of region A and C) to higher energies contributes to the flux. In practice, this transfer would either increase the number of escaping particles, enhance the anti-sunward portion of the halo (compared to the sunward portion), or both. Another refinement for better alignment with observations would be to account for diffusion creating a proper sunward deficit, rather than just an empty region, along with its associated negative flux contribution. The key idea is that the sum of all newly introduced flux contributions from diffusion (of the core) could ultimately result in a net increase in total flux, thereby providing an additional acceleration mechanism within the kinetic exospheric model.
@@
The observed suprathermal electrons, included in our existing kinetic exospheric model, constitute an important property of the strahl as they may have significant implications on the solar wind acceleration. According to Effenberger & Jeffrey (2021), spectral line analyzes seem to indicate that suprathermal populations are present in the corona. However, in situ data from PSP presented by Halekas et al. (2022) suggest that no significant strahl suprathermal electrons are observed near the Sun. This observation implies that Kappa electron VDFs may not be the most likely explanation for the fast solar wind. However, our results, consistent with the findings of Zouganelis et al. (2004), indicate that the key factor influencing the solar wind’s terminal velocity is not solely the suprathermal component, but rather the relative abundance of particles within the strahl compared to the deficit in the sunward direction.

### v3

@@ 3 Exospheric approach on diffusion
The main limitation of ordinary kinetic exospheric models of the solar wind is their inability to self-consistently define a halo population outside the regions of velocity space described in the previous section. To solve this problem, we introduced a halo population that is directly dependent on the number of particles in the classically filled regions of velocity space (A, B, and C). In this way, the radial evolution of the total density stays unchanged while the electrons are redistributed in velocity space so that a tenuous halo appears for greater perpendicular velocities.
Therefore, diffusion is modeled by defining a new parameter that aims to reduce the number of electrons in the A, B, and C regions and adding the same number of electrons to the initially empty region of Fig. 2 labeled D. This diffusion parameter is defined as
@@
so that when $r\rightarrow\infty$ ($\eta\rightarrow 0$) then $D_{r}\rightarrow\zeta$, with $\zeta$ being a value between 0 and 1 that modulates the relative number of diffused particles, compared to the others (of regions A, B, and C), at an infinite distance from the Sun. The choice that the diffusion profile evolves as $1-(r_{0}/r)^{6}$ is entirely arbitrary. It was selected to ensure that $D_{r}=0$ at $r=r_{0}$ while also preventing excessively rapid growth, thereby avoiding the diffusion coefficient reaching near its maximum value well before approaching 1 AU.
Furthermore, the density within the combined velocity space region of A, B, and C can be deduced by the addition of Eqs. (5), (25), and (30) of Pierrard & Lemaire (1996). This can be written as
@@
such that the density of electrons $n_{0}$ is equal to the total density of electrons at the exobase since
@@
Using the total density and the diffusion parameter, we can now deduce the density of the depleted region (combined region of A, B, and C), which corresponds to
@@
such that the total density remains unchanged since $n_{e}(r)=n_{e}^{\prime}(r)+n_{e,\textbf{D}}(r)$.
The electron flux is here defined using Eq. (26) of Pierrard & Lemaire (1996) by taking into account the diffusion. This corresponds to the only nonzero contribution of the flux associated with the escaping particles (region B) and written as
@@
From this, it follows that the total flux is $F_{e}^{\prime}(r)+F_{e,\textbf{D}}(r)=F_{e}(r)(1-D_{r})$, with $F_{e,\textbf{D}}(r)=0$ being the diffusion contribution to the flux and $F_{e}(r)$ being the total flux as if there were no diffusion. Here, the diffused electrons (from the depleted region) are scattered so that they do not contribute to the flux ($F_{e,\textbf{D}}(r)=0$). This provides a simple approach for formulating the effect of diffusion on the flux in our kinetic exospheric model. We note that the proton flux must be modified accordingly to maintain the zero-current condition that was used to find the solution for the electric potential.
Figure 2: Radial evolution of the electron VDF from the exobase (1.2 $R_{s}$ in our model) to 1 AU ($\sim$215 $R_{s}$) modeled by including diffusion for electrons. Region D represents the electron diffusion region where $\zeta=0.1$ was assumed.
To deduce the electron pressures of the D region, we first defined the pressure without any diffusion of electrons by adding Eqs. (7), (27), and (31) of Pierrard & Lemaire (1996) for the parallel component to obtain
@@
and adding Eqs. (8), (28), and (32) for the perpendicular component to obtain
@@
To determine the pressures in the diffusion region, we introduced the electron density $n_{e,\text{I}}$, obtained by integrating the electron VDF from Eq. (9) within the boundaries of the D region. This quantity corresponds to subtracting Eqs. (35) with (25) of Pierrard & Lemaire (1996), as expressed by
@@
In the same way, we defined the parallel and perpendicular pressure by subtracting Eq. (27) from Eq. (37) and Eq. (28) from Eq. (38), respectively, and multiplying them by the ratio $n_{e,\textbf{D}}(r)/n_{e,\text{I}}(r)$ to obtain
@@
The pressures are those corresponding to the D region of the electron VDF in Fig. 2, where $N_{e,\textbf{D}}$ is the normalization factor of the electron VDF (replacing $N_{e}$ in Eq. (9)) so that the density of this region is always equal to $n_{e,\textbf{D}}(r)$ as described by Eq. (31). This is inferred by noticing that
@@
This allowed us to then deduce the bulk velocity,
$v_{\text{sw}}(r)\equiv\frac{F_{e}^{\prime\prime}(r)}{n_{e}^{\prime\prime}(r)},$ (47)
and total temperatures,
@@ 4 Diffusive kinetic exospheric model: Results
To emphasize the impact of diffusion, we analyzed two scenarios: one that incorporates diffusion and another that does not. We adopted the same fundamental parameters of the exospheric model as those used for the illustrations in Figs. 1 and 2, as specified in Table 1, and we assumed the diffusion radial evolution described by Eq. (20) with $\zeta=0.1$ as the relative maximum amount of diffused particles. This $\zeta$ value was arbitrarily chosen to qualitatively analyze the effect of diffusion, as further improvements to the exospheric diffusion framework would be required to accurately constrain the radial evolution of diffused particles with observations. The radial evolution of the diffusion, electron temperatures, solar wind speed, and electron temperature anisotropy given by our kinetic exospheric model is shown in Fig. 3.
Figure 4: Radial evolution of the different electron population densities computed with the parameters from Table 1 and $\zeta=0.1$ in our modified kinetic exospheric model.
In Fig. 3a, the proportion of diffused particles increases with the radial distance, as expected from Eq. (20). The radial evolution of the electron temperatures, when affected by diffusion, shows a slower decrease in Fig. 3b. This behavior is explained by the fact that the electron VDF is essentially broadened by placing particles in the diffusion region. Fig. 3d shows a clear stabilization of the temperature anisotropy, which reaches a value of 1.65 at 215 $R_{s}$ when diffusion is considered. At 1 AU, observations show that the temperature anisotropy $T_{\parallel}/T_{\perp}$ is typically around 1 (that is, without anisotropy) for a slow wind of 300 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ and increases with the solar wind speed to 1.4 at 700 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ (Salem et al. 2023). In this respect, the naive introduction of diffusion in our model already brings the temperature anisotropy to a value fairly close to the observations in comparison to what was previously found with exospheric models.
The solar wind speed in Fig. 3c shows a slight decrease when diffusion is included. This is because the diffused particles are assumed to be redistributed without any preferential parallel direction. Therefore, these particles do not contribute to the bulk velocity but, being the only contribution to the flux of the solar wind in our model, do effectively diminish the number of particles within the strahl (essentially the B region). This in turn reduces the bulk velocity shown in Fig. 3c. As a consequence, because of the zero-current condition, the bulk velocity of protons must decrease by the same magnitude as that of the electrons.
@@ 5 Exospheric perspective on solar wind acceleration
As presented in Sect. 2, a key advantage of the kinetic exospheric model is its ability to self-consistently determine the electric field under the constraints of quasi-neutrality and zero-current conditions. Consequently, the ambipolar electric field arises directly from the greater influence of the gravitational field on protons compared to that of electrons. This resulting global electric potential contributes, at least partially, to the acceleration of the solar wind and is generally entirely responsible for the acceleration in the context of kinetic exospheric models (Berčič et al. 2021a, b).
@@
Following the kinetic exospheric approach, a signature of the ambipolar electric potential can be identified in the solar wind data. Specifically, it can be inferred from the cutoff velocity associated with the deficit in the sunward direction by analyzing the observed electron VDFs in the solar wind (Halekas et al. 2020; Berčič et al. 2021b, c; Halekas et al. 2021, 2022). Introducing the asymptotic wind speed, Halekas et al. (2022) binned the cutoff velocities of PSP between 13.3 and 50 $R_{s}$ (for encounters 3 to 11). In this way, the radial evolution of the electric potential associated with asymptotic speeds of 200–250 and 600–650 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ has been deduced and is represented in Fig. 5 for PSP slow solar wind (SSW) and PSP fast solar wind (FSW). An anticorrelation between the electric potential and the velocity of the solar wind can be clearly identified (Halekas et al. 2022).
@@
Figure 5: Radial evolution of the electric potential with respect to infinity. The solid lines are the electric potential retrieved by the exospheric model with the input parameters from Table 2. The corresponding solar wind terminal velocities of our model are approximately: 320 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for SSW1, 280 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for SSW2, 640 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for FSW1, and 520 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ for FSW2. The dashed lines correspond to the SSW (blue) and FSW (orange) limits fitted to PSP data between 13.3 and 50 $R_{s}$ (encounters 3 to 11) by Halekas et al. (2022). The dashed blue line is the fit of solar wind data with asymptotic speeds of 200–250 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$, while the dashed orange line corresponds to asymptotic speeds of 600–650 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$.
@@
For the kinetic exospheric model, the radial evolution of the self-consistent electric potential is shown in Fig. 5 for four different sets of parameters given by Table 2. The SSW1 solid black line is a SSW solution found using the same test parameters as in Table 1, which was introduced to illustrate the effect of diffusion. We note that the electric potential is independent of the diffusion, since it is found by using the quasi-neutrality with an unchanged total density. It is commonly understood that exospheric models predict a faster solar wind accompanied by a larger electric potential at any radial distance, to account for the enhanced acceleration of the solar wind. This contradiction with the observed anticorrelation mentioned above holds true only if we consider that the acceleration is entirely determined by the $\kappa$ parameter. As illustrated in Fig. 5, this is evident from the comparison between SSW1 and FSW1, where reducing $\kappa$ from 6 to 3 results in a corresponding increase in the terminal wind velocity from 320 to 640 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$. However, when considering the same change in $\kappa$ for SSW2 and FSW2, a carefully selected combination of exobase levels and initial temperatures (see Table 2) produces a crossing point between the two electric potentials (at around 13 $R_{s}$) while still maintaining the increase in terminal wind velocity from 280 (for SSW2) to 520 $\mathrm{k}\mathrm{m}\,\mathrm{s}^{-1}$ (for FSW2). These parameters were chosen because they approximate the observed behavior to some extent, producing a crossing point at low altitude while maintaining realistic values (as described in the next paragraph). However, a wide range of parameter sets can reproduce a crossing point, provided they meet two key conditions: (1) a sufficiently large difference between proton and electron temperatures (compared to the difference in kappa values) in fast wind relative to slow wind, and (2) a significantly higher exobase level for slow wind. In this way, the anticorrelation between the electric potential and the solar wind terminal velocity is retrieved above 13 $R_{s}$ with the kinetic exospheric model. This indicates that the kinetic exospheric model may be better suited than previously thought for explaining the FSW through the electric field, especially if the crossing point between the electric potentials of fast and slow winds occurs below 13 $R_{s}$. Although PSP observations do not appear to indicate the presence of such an intersection, or at least not above 13 $R_{s}$, the consistency of our modeled electric potentials compared with observations suggests that further investigation of the most recent PSP data is needed to completely rule out this possibility.
@@
This consideration is reinforced by the fact that the parameters found to reproduce the anticorrelation between the electric potential and the solar wind speed (see SSW2 and FSW2 of Table 2) are also consistent with the deduced coronal temperatures (David et al. 1998; Cranmer 2002). The FSW typically appears above coronal holes that are characterized by a higher proton temperature compared to that of electrons and a low density that lowers the exobase compared to other regions of the Sun’s surface (Pierrard et al. 2023), which is the case for FSW2. This is also supported by the anticorrelation between the strahl parallel temperature and the solar wind speed, consistently found by Helios (65–215 $R_{s}$) and PSP (13.3–50 $R_{s}$; Berčič et al. 2020; Halekas et al. 2022). Furthermore, as identified with SSW2, the SSW is commonly associated with denser and less active regions of the equatorial corona, where the exobase is therefore higher with a generally hotter electron temperature than the coronal hole conditions (David et al. 1998). As reported by Berčič et al. 2020, these findings suggest that the strahl parallel temperature preserves the electron coronal temperature. The exospheric approach is consistent with this idea by maintaining the strahl parallel temperature constant throughout the solar wind expansion, since no collisions nor wave-particle interactions are assumed.
@@
It is important to note that the implementation of the diffusion mechanism does not inherently require that particles in region D do not contribute to the flux. For that reason, an additional acceleration mechanism could be taken into account within the current diffusion framework by considering that the transfer of core particles (combined ballistic and trapped particles of regions A and C) to higher energies contributes to the flux. In practice, this transfer would increase the number of escaping particles and/or enhance the anti-sunward portion of the halo (compared to the sunward portion). Another refinement for better alignment with observations would be to account for diffusion creating a proper sunward deficit, rather than just an empty region, along with its associated negative flux contribution. The key idea is that the sum of all newly introduced flux contributions from diffusion (of the core) could ultimately result in a net increase in total flux, thereby providing an additional acceleration mechanism within the kinetic exospheric model.
@@
The observed suprathermal electrons, included in our existing kinetic exospheric model, constitute an important property of the strahl as they may have significant implications for the solar wind acceleration. According to Effenberger & Jeffrey (2021), spectral line analyzes seem to indicate that suprathermal populations are present in the corona. However, in situ data from PSP presented by Halekas et al. (2022) suggest that no significant strahl suprathermal electrons are observed near the Sun. This observation implies that Kappa electron VDFs may not be the most likely explanation for the FSW. However, our results, consistent with the findings of Zouganelis et al. (2004), indicate that the key factor influencing the solar wind’s terminal velocity is not solely the suprathermal component, but rather the relative abundance of particles within the strahl compared to the deficit in the sunward direction.